OpenStudy (okdutchman7):

Please help. WILL MEDAL! 2 ln 1 + 2 ln 5 = ln (x+3) @mathmale @hartnn @triciaal @sleepyjess @satellite73 @jim_thompson5910 @zepdrix @aaronq @Zale101

OpenStudy (marcelv):

Remembering that ln(1) = 0, we have then \[\ln(5^2) = \ln (x+3)\] that give us 25 = x+3, so x = 22

OpenStudy (okdutchman7):

@hartnn @jim_thompson5910 @satellite73 @sleepyjess @zepdrix

OpenStudy (photon336):

2Ln(1) = 1^2 = 1 Ln(1) = 0 2Ln(5) = ln(5^{2}) = ln(x+3) We can take both sides to the e^ because e^ln(x) = x e^ln(5^2) = e^ln(x+3) 5^2 = x+3

OpenStudy (triciaal):